OpenStudy (anonymous):

Find the smaller solution for 5x^2-51=9

OpenStudy (campbell_st):

add 51 to both sides 5x = 60 divide by 5 x^2 = 60/5 x^2 = 12 then take the square root of both sides \[x = \pm \sqrt{12}\] or simplifying \[x = \pm 2\sqrt{3}\] you need to decide which solution is smaller

OpenStudy (anonymous):

how do yo know which solution is smaller?

OpenStudy (campbell_st):

well it seems from what I've read on here... the negative is considered the smaller

OpenStudy (anonymous):

how do you know which is the negative?

OpenStudy (campbell_st):

well try \[x = -\sqrt{12}\] or \[x = -2\sqrt{3}\]

OpenStudy (anonymous):

so the answer would be -2

OpenStudy (campbell_st):

no the answer is \[x = -\sqrt{12} \] or \[x = -2\sqrt{3}\] you need the full answer including the square root